OpenStudy (anonymous):

9) The perimeter of a rectangle is 100 ft. and the length is 20 ft. What is the width?

OpenStudy (anonymous):

perimeter = Lx2 + Wx2. so do Lenght = 20x2 = 40. now do 100 - 40 = 60. then do 60/2 = 30. so width = 30 ft